40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 254 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  4759   Mon May 23 00:36:51 2011 kiwamuUpdateLSCDRMI trial : sucess

Eventually the DRMI was locked.

I was struggling to find a good signal port for SRC over the weekend and finally found AS55_I worked somehow. I used :

   REFL11_I --> PRC

   AS55_Q   --> MICH

   AS55_I    --> SRC

A configuration script was prepared such that someone can try this configuration by clicking a button on the C1IFO_CONFIGURE.adl screen.

I don't think this signal extraction scheme is the best, but now we can find better signal ports by shaking each DOF and looking at each signal port.

More details will be reported in the morning.

Quote:

I will try with POY55 that Koji prepared today.

 

  4758   Sat May 21 17:02:38 2011 KojiUpdateElectronicsAlberto's 11MHz was modified to POP55MHz

- Resonant at 55MHz. The transimpedance is 258Ohm. That is about half of REFL55 (don't know why).

- 11MHz&110MHz notch

- The 200MHz oscillation of MAX4106 was damped by the same recipe as REFL11.

POP55_transimpedance.pdf

 

  4757   Sat May 21 06:19:46 2011 kiwamuUpdateLSCDRMI trial : no luck

I will try with POY55 that Koji prepared today.

  4756   Fri May 20 11:37:44 2011 KojiUpdateElectronicsREFL11 fixed (REFL11 not working)

- Found the inductor which shunts the positive input of MAX4107 was not touching the ground.
This left the positive input level undetermined at DC. This was why MAX had been saturated.
The PCB has a cut, so it was surprising once the circuit worked.

- Resoldered the inductor to the ground. This made the circuit responding to the intensity-modulated beam.

- But the resonances and the notches were totally off, and the 200MHz oscillation has resurrected.

- Attached 40Ohm+22pF network between the neg-input of MAX and the gnd. This solved the oscillation.

- Made the tuning and the characterizations. The PD is on Kiwamu's desk and ready to go.

More to come later

Quote:
I took REFL11 out from the AS table for a health check because it wasn't working properly.

The symptoms were :

   - a big offset of ~ -3 V on the RF output. No RF signals.

   - The DC output seemed to be okay. It's been sensitive to light.

I did a quick check and confirmed that +/- 5V were correctly supplied to the op-amps.

It looks that the last stage (MAX4107) is saturated for some reasons. Need more inspections.

At the moment the REFL11 RFPD is on the bench of the Jenne laser.

 

  4755   Fri May 20 05:41:22 2011 SureshUpdateRF SystemDemod board measurements

The POP110 board which had the large Amplitude and Phase unbalances was examined today.  It turned out that there was some stray solder which had connected the Sum port of the PSCQ-2-120 splitter to its body (ground).  After I removed that the amplitude unbalance was 0.3dB however the phase was 105deg.  The phase reduced to 90 deg only if the power on the splitter is around 19 dBm.  So removed the AT1 (10dB attenuator) and the phase unbalance dropped to 91 deg.  However this is not a sustainable solution as the ERA-5 max output is about 19.5 dBm.

As this is a side band power monitor (and not a length sensing RFPD), we can make do with a poorer phase.  I will therefore replace the AT1 and adjust the residual phase with cable delay lines.  

Quote:

I measured the amplitude and phase imbalances of the demod boards which have been modified.  This is just a basic health check.  We hope to use the script that Kiwamu is developing for a more accurate test.  The script can also use these measurements as a sanity check.  POP110  requires some further attention. 

Demod_Board_measurements.png

 

The RF distribution box outputs corresponding to the demod board (eg. AS55_LO --> AS55_demod) were used as LO sources.  The RF signal was generated with a Marconi and held a kHz away from the LO frequency.  The amplitude and phase unbalance were measured with SR785.  The RF Power meter was used to check the LO power in each case.

 

 

 

  4754   Fri May 20 03:29:04 2011 kiwamuUpdateCDSBinary IO box on 1X5 : LEDs off

[Steve / Kiwamu]

 When Steve was working on the strain reliefs on 1X5 he found that some LEDs on the back side of the binary IO boxes were off.

There are 4 binary IO boxes and their power are directly supplied from Sorensens. According to the display of the Sorensens, the power are correctly generated.

Steve and I checked a picture of the boxes taken before he started working and we found it's been like this.

It might be just a problem of the LEDs or the fuses are blown, but anyway it needs an inspection.

Here is a picture of the back side of the boxes. You can see some LEDs are on and some are off.

DSC_3050_small.jpg

  4753   Fri May 20 03:01:17 2011 kiwamuUpdateLSClocking status

(PRMI locking)

Since REFL11 has gone I tried locking the PRMI with combination of REFL55 and AS55.

Without any pain the lock of PRMI was achieved successfully. AS55 was used to sense MICH and REFL55 was used for PRC.

 

(scripting)

Additionally I was modifying several scripts which are invoked from C1IFO_CONFIGURE.adl. Some details about the scripts will be uploaded on the wiki later.

An important thing is that now we are able to use the "restore" commands for the Y arm, X arm, Michelson and PRM locking.

The scripts will automatically acquire the lock of each DOF.  The image below is just a screen shot of the medm screen where you can call the scripts.

Screen_shot_2011-05-20_at_2.50.00.png

 

( Still to do)

   * PRM actuator response measurement

   * PRC noise budget

   * MICH-PRC actuator decoupling

  4752   Fri May 20 01:02:50 2011 loose connection hunterUpdateSUSloose connection on ETMY rack

The UL signal of the shadow sensor on ETMY went to zero this evening.

This was due to a loose connection on the cross connection board on the 1Y4 rack.

In order to make them tighten, a combination of stand-offs and screws were installed on the connectors. They won't be loose any more.

ETMY_loose.png

  4751   Thu May 19 19:41:17 2011 SureshUpdateRF SystemDAQ channel Assignments for RF PDs

[Kiwamu, Suresh]

In trying to keep the wiring of the LSC rack as neat as possible, we came up with the following channel assignments of the RF PD signals. 

PDI = PD Interface, The PD Interface D-type connectors (#1 to #12) are numbered:  Top -> Bottom and Left -> Right in ascending order. 

The Analog channels on the LSC Whitening Filter boards are numbered similarly, 1 to 32 in four sets.

1Y2_Rack_Layout.png

  4750   Thu May 19 17:53:03 2011 steveUpdateCDSAA filter box modified at 1X5

Quote:

[Steve, Joe]

Steve pulled the top AA filter box from 1X5 which handled some of the suspensions channels.  We turned off all the watchdogs before pulling it out, as well as recorded which cables were connected to which inputs.

The case  is undergoing a structural modification to have the ADC adapter card which previously was loosely connected via cables, securely attached to the case.

Steve still wants to do some cabling in the rack while the box is out, and will return it this afternoon once he has finished that.

 Job is done. Sus damping are back on. Cabling-strain reliefing are  not finished yet at 1X5 and 1X4

  4749   Thu May 19 16:46:20 2011 kiwamuUpdateLSCc1lsc model : input channels rearanged

According to Suresh's LSC rack design I rearranged the input channels of the c1lsc model such that the analog signals and the ADC channels are nicely matched.

Also I updated the c1lsc model in the svn with a help from Joe. The picture below is a screen shot of the input channels in the model file after I edited it.

c1lsc.png

  4748   Thu May 19 12:09:41 2011 josephbUpdateCDSAA filter box pulled from 1X5, optic suspensions currently off

[Steve, Joe]

Steve pulled the top AA filter box from 1X5 which handled some of the suspensions channels.  We turned off all the watchdogs before pulling it out, as well as recorded which cables were connected to which inputs.

The case  is undergoing a structural modification to have the ADC adapter card which previously was loosely connected via cables, securely attached to the case.

Steve still wants to do some cabling in the rack while the box is out, and will return it this afternoon once he has finished that.

  4747   Thu May 19 03:13:54 2011 kiwamuUpdateElectronicsREFL11 not working

I took REFL11 out from the AS table for a health check because it wasn't working properly.

The symptoms were :

   - a big offset of ~ -3 V on the RF output. No RF signals.

   - The DC output seemed to be okay. It's been sensitive to light.

I did a quick check and confirmed that +/- 5V were correctly supplied to the op-amps.

It looks that the last stage (MAX4107) is saturated for some reasons. Need more inspections.

At the moment the REFL11 RFPD is on the bench of the Jenne laser.

 

  4746   Thu May 19 00:23:44 2011 ranaUpdateCDSdiagonalization of MC input matrix

 I've moved all of my SOS peak fitting stuff into the scripts area so that Leo can make it better:

/cvs/cds/rtcds/caltech/c1/scripts/SUS/peakFit

findPeaks.m gets the data and makes the fitted spectra that I put in the previous entry.

findMatrix.m is the barely started script that ought to take the TF data and output the matrix to the MEDM screen.

  4745   Thu May 19 00:22:21 2011 ranaUpdateelogRestarted, Italian-style

Quote:

Aka, from a hotel in Pisa.

 Restarted Thu May 19 00:21:49 2011 to recover from Jenne's Italian terrorism.

  4744   Thu May 19 00:15:20 2011 JenneUpdateelogRestarted, Italian-style

Aka, from a hotel in Pisa.

  4743   Wed May 18 22:11:31 2011 ranaConfigurationLSCNew digital lockin added to LSC model/screen

Untitled.png

 This is OK....but, the input matrix should come from the same place as the regular input matrix: i.e. it should be just another row like CARM, DARM, etc. rather than have its own screen.

Also, I think a nice mod to all the matrices would be if the ORANGE triangle was only visible when there's a signal flowing through it.

  4742   Wed May 18 18:48:46 2011 kiwamuUpdateElectronicsidentification of RFPD interface cables

[Haixing / Kiwamu]

 As a part of the Wednesday's cabling work, we spent some times for identifying the RFPD interface cables.

The RFPD interface cables are made of a 15 pin flat cable, containing DC power conductors for the RFPDs and the DC signal path.

The list below is the status of the interface cables.

 

- - - - RFPD name, (cable status) - - - -

- REFL11 (identified and labeling done)

- REFL33 (identified and labeling done)

- REFL55 (identified and labeling done)

- REFL165 (no cable found)

- AS55 (identified and labeling done)

- AS165 (identified and labeling done)

- POP22/110 (identified and labeling done)

- POX11 (identified and labeling done)

- POY11 (identified and labeling done)

- POY55 (identified and labeling done)

We still have two cables which are not yet identified. Their heads are around the LSC rack and labeled 'unidentified'

  4741   Wed May 18 18:33:46 2011 AidanUpdateelogrestarted elog with script
  4740   Wed May 18 17:06:39 2011 SureshUpdateRF SystemDC power supplies for the RF generation box in place

 

I have checked the voltages on the connector.  They are okay and I have plugged in the Sorensen power into the RF Source.  The ground reference for the Sorensens comes from the 1X2 Rack ground reference lines on the south side of the rack. 

I looked for the OMC ground reference. Could not find one on either of the the OMC half racks.

Quote:

[Koji, Steve]

DC power supplies for the RF generation box are now in place. They are the top two of the 6 Sorensens in the OMC short rack next to 1X2.

We made the connections as we did for the RF distribution box, the power supplies labele, and the cables strain-relieved.

The power supply is not yet connected to the actual RF generation box. This should be done by Suresh or someone with the supervision of him.

Note:
We have two +18V supply on the short OMC rack, in total. One is for the RF source, the other is for the OMC PZTs, whitening, etc.
This is to avoid unnecessary ground loop
although the grounding situation of the OMC side is not known to me.

 

  4739   Wed May 18 16:52:23 2011 SureshUpdateRF SystemCables for AS11 PD are in place

[Larisa, Suresh]

All the cables needed for the AS11 PD are in place... the heliax cable runs from the AS table to the PSL rack.  The LO and RF cables to demod board as well as the I and Q cables into the LSC Whitening board are connected.

The cables get rather densely packed when the LSC Whitening filter sits between the PD Interface Board and the LSC AA filter board.  This makes it difficult to access the SMA connectors on the LSC whitening filter.  So we shifted the LSC Whitening and AA Filter boards one slot to the right.  The LSC rack looks like this just now.  We have also shifted the binary cables at the back of the Eurocart by one slot so the same cables are associated with the cards.

1Y2_Rack_Layout_short-term.png

 

 

  4738   Wed May 18 15:54:50 2011 KojiUpdateRF SystemDC power supplies for the RF generation box in place

[Koji, Steve]

DC power supplies for the RF generation box are now in place. They are the top two of the 6 Sorensens in the OMC short rack next to 1X2.

We made the connections as we did for the RF distribution box, the power supplies labele, and the cables strain-relieved.

The power supply is not yet connected to the actual RF generation box. This should be done by Suresh or someone with the supervision of him.

Note:
We have two +18V supply on the short OMC rack, in total. One is for the RF source, the other is for the OMC PZTs, whitening, etc.
This is to avoid unnecessary ground loop
although the grounding situation of the OMC side is not known to me.

  4737   Wed May 18 07:18:15 2011 SureshUpdateRF SystemPOY55_Demod board Hardware change completed

The ~350 Hz noted in the elog below was traced to an RF modulation of the 11 MHz sideband.  This modulation was set up in the Marconi which is currently supplying the 11 MHz local oscillator signal to the RF source.  lt was used during the MC length study completed last week by Valera and Ryan.  The frequency measured was 322 Hz.

As we do not require this any longer, I have switched off this modulation.

 

 

Quote:

The Demod board with S. No. 022 (being used earlier as REFL11) has been modified.  It now has SCLF-65 as its input LP filter on the PD input line and a PQW-2-90 power splitter.  The unit functioning okay (I and Q signals are 90 deg apart.

The loss of Q output was traced to a possible loose solder joint and we now have both the I and Q signals after resoldering all components in the vicinity of U7 (Ref Schematic of D990511)

There is a strong oscillation around 350Hz present on I and Q signals of both REFL55_Demod and POY55_Demod.  Don't know the source. 

We have run out of power splitters to continue with the Demod board modification. We do not currently have an AS11_Demod board.  All the others are in place and ready for the I<->Q phase angle measurement.

In summary we now have the following Demod boards in place:

[ REFL11, POY11, REFL55, AS55, POY55, POY22, POY110]_Demod

 

 

  4736   Wed May 18 07:13:00 2011 SureshUpdateRF SystemDemod board measurements

I measured the amplitude and phase imbalances of the demod boards which have been modified.  This is just a basic health check.  We hope to use the script that Kiwamu is developing for a more accurate test.  The script can also use these measurements as a sanity check.  POP110  requires some further attention. 

Demod_Board_measurements.png

 

The RF distribution box outputs corresponding to the demod board (eg. AS55_LO --> AS55_demod) were used as LO sources.  The RF signal was generated with a Marconi and held a kHz away from the LO frequency.  The amplitude and phase unbalance were measured with SR785.  The RF Power meter was used to check the LO power in each case.

 

 

  4735   Tue May 17 22:50:00 2011 JamieConfigurationLSCNew digital lockin added to LSC model/screen

I added a lockin to the LSC model, and added it's corresponding control to the LSC screen.  Here's is what I added to the LSC model:

lsc-mdl-lockin.png

On the left are the froms from the normalized RF PD outputs. Those go into a matrix, whose single row goes into the lockin signal input. The clock output from the lockin goes into the LSC output matrix (last row).

Here is what I added to the LSC master screen:

lsc-screen-lockin.png

I also modified the lockin overview screen:

lsc-lockin-screen.png

I think this new screen looks a lot cleaner.  Maybe we could start using this one as a template for lockin screens.

  4734   Tue May 17 19:38:32 2011 kiwamuUpdateSUSwrong connection on 1X5

Today Steve was working around the 1X5 rack to strain relief the cable jungles and the jungle is now getting less jungle.

During the work he disconnected and reconnected some cables.

So for a doublecheck I checked all the suspensions to see if the suspensions are still healthy or not.

Aha, then I found a mistake.

 

See the pictures below. It's a very subtle difference. This wrong connection prevented MC1 and MC3 from damping.

wrong.png correct.png

  4733   Tue May 17 18:09:13 2011 Jamie, KiwamuConfigurationCDSc1sus and c1auxey crashed, rebooted

c1sus and c1auxey crashed, required hard reboot

For some reason, we found that c1sus and c1auxey were completely unresponsive.  We went out and gave them a hard reset, which brought them back up with no problems.

This appears to be related to a very similar problem report by Kiwamu just a couple of days ago, where c1lsc crashed after editing the C1LSC.ini and restarting the daqd process, which is exactly what I just did (see my previous log).  What could be causing this?

  4732   Tue May 17 17:01:22 2011 jamieConfigurationCDSUpdate LSC channels from _DAQ to _DQ

As of RCG version 2.1, recorded channels use the suffix "_DQ", instead of "_DAQ".  I just rebuilt and installed the c1lsc model, which changed the channel names, therefore hosing the old daq channel ini file.  Here's what I did, and how I fixed it:

$ ssh c1lsc
$ cd /opt/rtcds/caltech/c1/core/trunk
$ make c1lsc
$ make install-c1lsc
$ cd /opt/rtcds/caltech/c1/scripts
$ ./startc1lsc
$ cd /opt/rtcds/caltech/c1/chans/daq
$ cat archive/C1LSC_110517_152411.ini | sed "s/_DAQ/_DQ/g" >C1LSC.ini
$ telnet fb 8087
daqd> shutdown

 

  4731   Tue May 17 16:53:22 2011 kiwamuUpdateSUSoplev sign was wrong on ETMY

Last night I found that the sign of the oplev control of PITCH on ETMY was wrong. I flipped it to the correct sign.

We've been locking the Y arm by feeding a signal back to ITMY  because pushing ETMY somehow made the lock unstable in the angular motion.

After the correction of the oplev contol sign, I was able to keep the lock robustly by pushing ETMY.

  4730   Tue May 17 11:45:20 2011 JamieConfigurationCDSpurged non-c1 site files from rtcds checkout of cds_user_apps
I purged all of the working copy checkouts of site files for all sites that are not c1 from the rtcds cds_user_apps working directory (/opt/rtcds/caltech/c1/userapps/trunk). I first checked that there were no outstanding changes, and then did the following (in bash):
cd /opt/rtcds/caltech/c1/userapps
rm -rf trunk
svn update --depth=files trunk
svn update --depth=empty trunk/{CDS,ISI,ISC,PSL,SUS}
svn update trunk/{CDS,ISI,ISC,PSL,SUS}/{c1,common}
  4729   Tue May 17 01:05:56 2011 kiwamuUpdateLSCAsymmetry measurement prep : recentering works

I re-centered beams on several PDs and a camera including :

  AS55, ETMY_QPD, TRY and ETMYT_CCD.

 

The most important one was AS55.

When I was locking each arm I found that the error signal from AS55 was very coupled to the angular motion of the arms.

I checked the beam on the AS55 RFPD and found the beam on the edge of the photo diode. This is possibly because Valera and I had been touching the input beam alignment.

At that time the DC signal from AS55 without aligning PRM and SRM was about 5 mV.

Adjusting the beam position by a steering mirror brought the DC signal up to 20 mV.

Then the lock of each arm became more stable.

  4728   Mon May 16 17:11:28 2011 steveUpdateElectronicsstrain relieved 1X1

1X1 is strain relieved in the back. I will use similar approach on the rest of the racks .

  4727   Mon May 16 15:46:05 2011 KojiUpdateGeneraloptical layout viewable on the online version of autocad

I have made my own AutoCAD WS account and put the latest 40m layout.
AutoCAD WS is a free cloud service which enable us to browse/edit the DWG files.

You can view/play/print it from the following link even without making the account.

https://www.autocadws.com/main/publish?link=RVFmNEtOd3EzNVFtRXcx

If you make your account you can actually save it although the editing capability is somewhat limited.

Notes:
- It needs the Flash plug-in and Internet connection.
- I think Safari is the most stable platform on Mac. (i.e. I had some malfunctions with Firefox and Chrome)

  4726   Mon May 16 11:47:59 2011 kiwamuUpdateASSc1ass update part II

The medm screen for c1ass started being modified to be more user-friendly.

The modification is still ongoing, but the goal is to make a screen which anyone can easily understand and play with.

 

Still to do : ( need a volunteer )

- Modification of the screens

- Commission the input beam and X-arm servos

- Make scripts for X-arm

- Measure the PZT mirrors' matrix for the translation and angle

c1ass.png

Quote from #4709

Here the status of the dither alignment or c1ass:

Still to do:

- Commission the input beam and X-arm servos

- Make scripts for X-arm

 

  4725   Mon May 16 11:14:36 2011 ranaUpdateIOOwiring diagram for IP-POS

Just FYI, the 3113 is a 12-bit ADC, so we won't get very good resolution out of these. We should get one of those purple boxes.

Also, I corresponded some with Dave Barker. The 40m Wiki at LHO is down because of disk errors. He's working on it and will let us know. But suggests that we move it to Caltech since he'll turn the box off at the end of the year.

Mon May 16 13:57:49 2011    Wiki is back up.

  4724   Mon May 16 10:05:02 2011 kiwamuUpdatePhotosRe:ETMY optical bench

You are right. We should change or rotate the mirror mount.

Actually when Suresh and I were putting the mirror we rotated the mount  by 90 deg such that the fat side of the mount is at left had side.

It was because the fat side had been clipping the oplev beam when the fat side is at right.

At that moment we were blocking the green beam to only see the faint IR beam with a sensor card, so we haven't checked the green beam.

Anyway the mount is apparently not good for the green beam.

Quote from #4723

I didn't notice it the other day when I was working on putting in the trans QPD, but do we need to switch the mirror mount for the first turning mirror of the IR trans beam, which the green transmits through to go into the cavity?  It seems like we've set ourselves up for potential clipping.

 

  4723   Sun May 15 21:27:51 2011 JenneUpdatePhotosETMY optical bench

I didn't notice it the other day when I was working on putting in the trans QPD, but do we need to switch the mirror mount for the first turning mirror of the IR trans beam, which the green transmits through to go into the cavity?  It seems like we've set ourselves up for potential clipping.

Quote:

Just for a record. This is the latest picture of the ETMY optical bench.

I will upload this picture on the wiki after the wiki gets up.

ETMY_ss.jpg

 

  4722   Sun May 15 19:55:15 2011 kiwamuUpdatePhotosETMY optical bench

Just for a record. This is the latest picture of the ETMY optical bench.

I will upload this picture on the wiki after the wiki gets up.

ETMY_ss.jpg

  4721   Sun May 15 19:10:12 2011 kiwamuUpdateLSCCalibration of actuators : BS, ITMX and ITMY

The AC response of the actuators on BS, ITMX and ITMY were re-measured by another technique.

Last time I estimated them by measuring the open-loop transfer functions, but this time the responses were measured in a more direct way.

The measured AC responses (60 Hz - 200 Hz) are :

      BS   = 1.643e-98 / f2  [m/counts] (corrected based on the plot below - Manasa)

     ITMX = 3.568e-9 / f2 [m/counts]

     ITMY = 3.542e-9 / f2 [m/counts]

Next : measurement of the PRM actuator response


(The technique) 

 This time a technique that Rana told me a week ago was used.

This technique allows us to directly measure the response of an actuator at high frequency without any loop corrections.

First of all, MICH has to be locked to keep MICH within the linear range of the error signal. So now MICH is a linear sensor to the mirror motions.

In the MICH control a steep low pass filter should be inserted in order to avoid unwanted effects from the control loop at the high frequencies.

For example I put a low pass filter composed of an elliptical filter whose cut-off frequency is at 50 Hz such that the control loop doesn't push the mirrors above the cut-off frequency.

Hence the error signal of MICH above 50 Hz directly corresponds to the motion of the mirrors including BS, ITMX and ITMY.

Taking a transfer function from an actuator to the MICH error signal directly gives the actuator response.

In my measurements MICH was locked by feeding the signal back to BS. The plot below is the expected open-loop transfer function for the MICH control.

oltf.png

You can see that the open loop TF suddenly drops above 50 Hz. The UGF was at about 20 Hz, confirmed by looking at the loop oscillation on DTT.

 

(Measurement)

 In the technique the error signal has to be calibrated to [m]. This time AS55_Q was used and calibrated based on a peak-to-peak measurement.

The peak to peak value in the MICH error signal was 8 counts, which corresponds to the sensor efficiency of 4.72e+07 [counts/m].

Then I took transfer functions from each suspension (i.e. C1:SUS-XXX_LSC_EXC) to the error signal at AS55_Q over a frequency range from 60 Hz to 200Hz.

For the transfer function measurements I ran the swept sine on DTT to get the data. Note that the PD whitening filters were on.

The plot below is the results of the measurements together with the fitting lines.

calib_actuators.png

In the fitting I excluded the data pints at 60 Hz, because their coherence was low due to the power line noise.

  4720   Sun May 15 14:01:56 2011 kiwamuUpdateIOOwiring diagram for IP-POS

Here is a wiring diagram which shows how IP-POS (new official name is IB_POS) is connected.

Another thing we have to remember is : at some point we will also connect some more QPDs (e.g. POX, POY, AS, REFL and POP) in the same way.

They will also be acquired by the same slow machine : c1iscaux.

ip_pos.png

Quote from #4704

IP_POS is back. 

Hopefully I will make a simple wiring diagram such that we will never forget the connections.

  4719   Sun May 15 12:42:29 2011 kiwamuUpdateSUSf2p ratio adjutment done for all the suspensions

The f2p adjustment for all the suspensions are done (except for MC1,2,3)

  4718   Sun May 15 03:58:19 2011 ranaUpdateCDSdiagonalization of MC input matrix

There has been some input matrix diagonalization in the past by Yuta and Kiwamu, but I find the automation to be not totally satisfactory.

It would be better if we could automatically fit the data to find the Suspended optic eigenfrequencies and then use that to get the matrix. So I wrote a peak fitter to get the matrix.

It gets the data from mafalda with NDS2, then it makes the PSDs, and then starts with some initial guesses (based on looking at the plots) and them uses fminsearch to get the peak frequencies and Q's.

Using the output of this, we can use Yuta's method and take the passive transfer functions with the free swing data (from April 30, so we got do do it quick) to get the input matrix.

 

Doing the SUS input matrix is nice for having good damping (as long as we remember to include SIDE), but my motivation is to produce a good null stream from the 4 face sensors so that we can estimate the sensor noises at all times.

  4717   Sat May 14 14:50:21 2011 KojiUpdateLSC1Y2 5V Blown Fuse found -> Fixed

Incidentally, a blown fuse on 5V line at 1Y2 rack was found during the intallation of Sorensens.
The fuse (5A 125V) has been replaced and fixed.

When I plugged the fuse in, I heard some sound like relays were switched. Are there any relays in the LSC rack?

It was a 9th fuse from the top as seen in the picture.

  4716   Sat May 14 14:12:16 2011 KojiUpdateRF SystemDC power supply on RF distribution box has been replaced.

Key points of the power supply installation

  • We followed the grounding configuration for KEPCO except for the signal ground connection
  • AC power supply has been obtained from the local power strip. This also provides chassis earthing (for safety)
  • The chassis is connected to the shieldin of the DC supply cable. The other end should be isolated.
  • The low voltage side of Sorensen's DC outputs are connected in order to share the same reference  level.
  • The ground level is provided from the cross connect. The cable is connected between the cross connect ground to the sorencen.
    Unlike the KEPCO case, this cable does not have the current return, but just is to define the voltage level of those Sorensens.
  • New AC&DC cables have been nicely strain-relieved.

Quote:

[Steve, Koji, Suresh]

   We shifted two Sorensen power supplies from the Auxiliary rack next to 1X2 to 1Y2.  And have installed them there (pic below).  The local ground reference was picked up from the racks ground reference.  A shielded cable with two twisted pairs was used to make a new power cable for the RF rack.  Since we are using three of the four conductors (+18,+28 and ground), one of them is not connected to anything.  This situation can be improved in a future iteration when, for example, we might wish to relocate the Sorensens to a different rack.

   We are still working on changing the power supply to the RF source.  Will complete this early next week

 

  4715   Fri May 13 23:04:58 2011 SureshUpdateRF SystemDC power supply on RF distribution box has been replaced.

[Steve, Koji, Suresh]

   We shifted two Sorensen power supplies from the Auxiliary rack next to 1X2 to 1Y2.  And have installed them there (pic below).  The local ground reference was picked up from the racks ground reference.  A shielded cable with two twisted pairs was used to make a new power cable for the RF rack.  Since we are using three of the four conductors (+18,+28 and ground), one of them is not connected to anything.  This situation can be improved in a future iteration when, for example, we might wish to relocate the Sorensens to a different rack.

   We are still working on changing the power supply to the RF source.  Will complete this early next week

P5130120.JPG

  4714   Fri May 13 22:45:37 2011 SureshUpdateRF SystemThe full set of 8 Demod boards is ready for testing

We have Completed the hardware changes to the full set of 8 demod boards.  The last one completed today is AS11.  I have collected the info on all the demod boards available so far in the table below.  As we measure the actual phase and amplitude unbalance we will expand this table to include new info.

The set of 8 demod boards
Demod Board S. No. Power Splitter Frequency range of splitter (MHz) Phase unbalance from datasheet (deg) Amplitude Unbalance from datasheet (dB)
AS11 121 PSCQ-2-51W 5 to 50 87.49 0.1
REFL11 021 PSCQ-2-51W 5 to 50 87.49 0.1
POY11 119 PSCQ-2-32 3.2 to 32 87.58 0.05
AS55 029 PSCQ-2-51W 5 to 50 no info no info
REFL55 118 PQW-2-90 30 to 90 90.14 0.92
POY11 119 PSCQ-2-32 3.2 to 32 87.58 0.05
POY22 020 PSCQ-2-32 3.2 to 32 90.26 0.02
POY110 120 PSCQ--120 80 to 120 90.88 0.58

 

  4713   Fri May 13 17:20:48 2011 Leo SingerConfigurationSUSTuned bounce and roll mode of ETMY suspension

I tuned the bounce and roll mode bandstops for ETMY, although it was difficult for me to tell if there was improvement with the bandstops on relative to the bandstops off because it seemed like the bounce and roll modes were being excited intermittently.  I'll take spectra with the filters both on and off during an evening next week.

  4712   Fri May 13 14:54:20 2011 Leo SingerUpdateComputersDiaggui fixed on pianosa

I fixed diaggui on pianosa.  Previously, it was not able to start because it depended on libreadline5, whereas Ubuntu distributed libreadline6.  Now pianosa has both libreadline5 and libreadline6, so diaggui works.

  4711   Fri May 13 01:51:56 2011 SureshUpdateRF SystemRF Status update

I have posted the attached RF status update and 1Y2 rack layout to the svn.

  4710   Fri May 13 01:42:35 2011 kiwamuUpdateLSCready for Schnupp asymmetry measurement

[Valera / Kiwamu]

 We are able to lock each arm smoothly.  It is ready for the Schnupp asymmetry measurement.

( to be done )

 - Manual D-phase adjustment of the AS55 channel.

 - A script to adjust the D-phase.

 - Trigger logic for the boost filters.

 - Modification of some old alignment scripts to adopt them to the new LSC channels

ELOG V3.1.3-